OpenStudy (anonymous):

Determine whether the infinite geometric series is convergent or divergent. If it is convergent, find its sum. (If the series is divergent, enter DIVERGENT.) (1)-(1/6)+(1/36)-(1/216)+...

OpenStudy (anonymous):

absolutely converges

OpenStudy (anonymous):

How do I find the sum? I have no idea.

OpenStudy (anonymous):

r=- 1/6

OpenStudy (anonymous):

1 ------------- 1 - -1/6

OpenStudy (anonymous):

that doesn't work

OpenStudy (anonymous):

\[\huge {\color{DarkRed} {It\: doesent}}\]

OpenStudy (anonymous):

1/ 7/6 = 6/7

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

Everyone please HELP!

OpenStudy (anonymous):

The geometric series given is \[\sum_{n=0}^{\infty} (\frac{-1}{6})^n\] Which clearly converges since \(|-\frac{1}{6}|<1\).

OpenStudy (anonymous):

I'm sure you have the formula and can find the sum.

OpenStudy (anonymous):

So is it -1/6? If it is, my hw program says it's wrong.

OpenStudy (anonymous):

The answer is 6/7

OpenStudy (anonymous):

What's \(-\frac{1}{6}?\) The sum is not -1/6.

OpenStudy (anonymous):

it marked it wrong though!

OpenStudy (anonymous):

did you put 6/7?

OpenStudy (anonymous):

YEP!

OpenStudy (anonymous):

then I don't know what to say, 6/7 is correct answer

OpenStudy (anonymous):

im with imran. that is really the correct answer.

OpenStudy (anonymous):

Huh... I trust you guys. I guess there might be a glitch in the program then.

OpenStudy (anonymous):

GOT IT!!! Thank you :D

OpenStudy (anonymous):

Can you help me with this one? The tenth term of an arithmetic sequence is (57/2), and the second term is (9/2). Find the first term.

OpenStudy (anonymous):

9/2 + d (8)=57/2 find d

OpenStudy (anonymous):

subtract d from 9/2

OpenStudy (anonymous):

So D=3?